### v1

@@ Abstract
Simulating low-likelihood high-impact extreme weather events in a warming world is a significant and challenging task for current ensemble forecasting systems. While these systems presently use up to 100 members, larger ensembles could enrich the sampling of internal variability. They may capture the long tails associated with climate hazards better than traditional ensemble sizes. Due to computational constraints, it is infeasible to generate huge ensembles (comprised of 1,000-10,000 members) with traditional, physics-based numerical models. In this two-part paper, we replace traditional numerical simulations with machine learning (ML) to generate hindcasts of huge ensembles. In Part I, we construct an ensemble weather forecasting system based on Spherical Fourier Neural Operators (SFNO), and we discuss important design decisions for constructing such an ensemble. The ensemble represents model uncertainty through perturbed-parameter techniques, and it represents initial condition uncertainty through bred vectors, which sample the fastest growing modes of the forecast. Using the European Centre for Medium-Range Weather Forecasts Integrated Forecasting System (IFS) as a baseline, we develop an evaluation pipeline composed of mean, spectral, and extreme diagnostics. Using large-scale, distributed SFNOs with 1.1 billion learned parameters, we achieve calibrated probabilistic forecasts. As the trajectories of the individual members diverge, the ML ensemble mean spectra degrade with lead time, consistent with physical expectations. However, the individual ensemble members’ spectra stay constant with lead time. Therefore, these members simulate realistic weather states during the rollout, and the ML ensemble thus passes a crucial spectral test in the literature. The IFS and ML ensembles have similar Extreme Forecast Indices, and we show that the ML extreme weather forecasts are reliable and discriminating. hese diagnostics ensure that the ensemble can reliably simulate the time evolution of the atmosphere, including low likelihood high-impact extremes. In Part II, we generate a huge ensemble initialized each day in summer 2023, and we characterize the statistics of extremes.
@@ 1 Introduction
Recent low-likelihood, high-impact events (LLHIs) have raised important and unanswered questions about the recurrence times of these events and their relationship to anthropogenic climate change. For example, Hurricane Harvey in 2017 and the Summer 2021 heatwave in the Pacific Northwest (PNW) are two high-impact events with no modern analog. Several threads motivate research on LLHIs. First, the IPCC states that "the future occurrence of LLHI events linked to climate extremes is generally associated with low confidence" (Seneviratne et al., 2021, pp. 1536). Second, the occurrence of recent LLHIs, like the Summer 2021 PNW heatwave, reveals that the ability to characterize, let alone anticipate, such events is currently incomplete (Bercos-Hickey et al., 2022; Zhang et al., 2024; Liu et al., 2024).
@@ 2.1 Selecting an emulator
We choose SFNO because its spherical design is well-suited for problems in earth science, and the architecture includes scalable model parallel implementations (Bonev et al., 2023; Kurth et al., 2023). Since SFNO can be split across multiple GPUs during training, we can train large SFNOs and assess the effect of the SFNO size on the ensemble diversion. Additionally, for the same model size and hyperparameters, the compute complexity of SFNO scales linearly with horizontal resolution. This characteristic ensures that future versions of the emulator and the ensemble can efficiently handle datasets with high horizontal resolution, such as those from kilometer-scale climate simulations. A possible caveat is that higher resolution inputs may require more learnable parameters to maintain skillful prediction. In practice, this could still lead to super-linear compute complexity, if the total size of SFNO itself must also be increased. Future research is necessary to test SFNO on kilometer-scale datasets and to compare the practical compute costs with physics-based simulations.
@@ 3 Ensemble Diagnostics
Ultimately, with SFNO-BVMC, we hope to analyze the statistics of LLHIs. This requires a calibrated ensemble with reliable probabilistic forecasts. SFNO-BVMC is a novel way to create ensemble forecasts from deterministic ML models. Therefore, in the following section, we present a diagnostics pipeline to evaluate the SFNO-BVMC ensemble and compare it to the IFS ensemble. We first evaluate SFNO-BVMC using diagnostics that evaluate overall performance. Next, we assess SFNO-BVMC’s control, perturbed, and ensemble mean spectra. Finally, we present diagnostics specifically focused on extreme weather forecasts. We open-source the code for these diagnostics (see Data Availability section), and we hope that it can be used to guide future ML model development. For a fair comparison for all diagnostics, we validate IFS against ECMWF’s operational analysis and SFNO-BVMC against ERA5. All diagnostics show SFNO-BVMC results with 58 members, and IFS ENS results with 50 members. SFNO-BVMC has 58 members: 29 checkpoints and 1 bred vector per checkpoint (added to and subtracted from the initial condition). Because of the use of 29 checkpoints and centered bred vector perturbations, SFNO-BVMC cannot be evaluated with an ensemble size smaller than fifty-eight members. However, the difference in the metrics due to different ensemble size would be sufficiently small that the diagnostics still allow for fair evaluation.
@@ Acknowledgements.
There are four key components for the code and data used in this manuscript. For reproducibility, the training code is stored with a Zenodo DOI at https://zenodo.org/records/13137296, and the earth2mip code is stored with a Zenodo DOI at https://zenodo.org/records/13141566.
  1. 1.
    Trained SFNO Models: We open-source the learned weights of our models at https://portal.nersc.gov/cfs/m4416/earth2mip_prod_registry/. Each model is trained with a different initial seed: therefore, the name of each trained model end with "seed[SEED]," where SEED corresponds to the Pytorch seed used for the weight initialization. Each model also includes a ‘config.json‘ file which specifies the configuration parameters used for training each SFNO checkpoint.
  2. 2.
    Ensemble Inference with SFNO models: NVIDIA’s earth2mip repository can be used for ensemble inference with the SFNO checkpoints above. We use the version of earth2mip for inference at https://github.com/ankurmahesh/earth2mip-fork/tree/HENS. This version of earth2mip includes our implementation of bred vectors (earth2mip-fork/earth2mip/ensemble_utils.py) and ensemble inference (earth2mip-fork/earth2mip/inference_ensemble.py). The data portal gateway at https://portal.nersc.gov/cfs/m4416/ includes the other external dependency files necessary for ensemble inference with earth2mip.
  3. 3.
    Code to train the SFNO checkpoints: For training the SFNO checkpoints, we use modulus-makani v0.1.0. It is available at https://github.com/NVIDIA/modulus-makani, and we include some minor modifications to support training on Perlmutter at https://zenodo.org/records/13137296. This represents a frozen version of the code we used for training SFNO.
  4. 4.
    Code to score the SFNO checkpoints: https://github.com/ankurmahesh/earth2mip-fork/tree/scoring. We use a different branch of our fork of the earth2mip repository for scoring, including overall diagnostics and extreme diagnostics. After completing ensemble inference, we added more code to this branch for scoring. We separate point 4 to ensure there is a frozen version of the ensemble inference code available. The extreme diagnostics are scored for IFS vs operational analysis and SFNO-BVMC vs ERA5. The scoring code is in earth2mip-fork/earth2mip/extreme_scoring_utils.py and earth2mip-fork/earth2mip/score_ensemble_outputs.py, and earth2mip-fork/earth2mip/score_extreme_diagnostics.py. This branch includes additional scoring code compared to the branch shared in Point (3), but the ensemble inference code is the same.

### v2

@@ Abstract
Simulating low-likelihood high-impact extreme weather events in a warming world is a significant and challenging task for current ensemble forecasting systems. While these systems presently use up to 100 members, larger ensembles could enrich the sampling of internal variability. They may capture the long tails associated with climate hazards better than traditional ensemble sizes. Due to computational constraints, it is infeasible to generate huge ensembles (comprised of 1,000-10,000 members) with traditional, physics-based numerical models. In this two-part paper, we replace traditional numerical simulations with machine learning (ML) to generate hindcasts of huge ensembles. In Part I, we construct an ensemble weather forecasting system based on Spherical Fourier Neural Operators (SFNO), and we discuss important design decisions for constructing such an ensemble. The ensemble represents model uncertainty through perturbed-parameter techniques, and it represents initial condition uncertainty through bred vectors, which sample the fastest growing modes of the forecast. Using the European Centre for Medium-Range Weather Forecasts Integrated Forecasting System (IFS) as a baseline, we develop an evaluation pipeline composed of mean, spectral, and extreme diagnostics. Using large-scale, distributed SFNOs with 1.1 billion learned parameters, we achieve calibrated probabilistic forecasts. As the trajectories of the individual members diverge, the ML ensemble mean spectra degrade with lead time, consistent with physical expectations. However, the individual ensemble members’ spectra stay constant with lead time. Therefore, these members simulate realistic weather states during the rollout, and the ML ensemble thus passes a crucial spectral test in the literature. The IFS and ML ensembles have similar Extreme Forecast Indices, and we show that the ML extreme weather forecasts are reliable and discriminating. hese diagnosticsy ensure that the ensemble can reliably simulate the time evolution of the atmosphere, including low likelihood high-impact extremes. In Part II, we generate a huge ensemble initialized each day in summer 2023, and we characterize the statistics of extremes.
@@ 1 Introduction
Recent low-likelihood, high-impact events (LLHIs) have raised important and unanswered questions about the drivers of these events and their relationship to anthropogenic climate change. For example, Hurricane Harvey in 2017 and the Summer 2021 heatwave in the Pacific Northwest (PNW) are two high-impact events with no modern analog. Several threads motivate research on LLHIs. First, the IPCC states that "the future occurrence of LLHI events linked to climate extremes is generally associated with low confidence" (Seneviratne et al., 2021, pp. 1536). Second, the occurrence of recent LLHIs, such as as the Summer 2021 PNW heatwave, reveals that the ability to characterize, let alone anticipate, such events is currently incomplete (Bercos-Hickey et al., 2022; Zhang et al., 2024; Liu et al., 2024).
@@ 2.1 Selecting an emulator
We choose SFNO because its spherical design is well-suited for problems in earth science, and the architecture includes scalable model parallel implementations (Bonev et al., 2023; Kurth et al., 2023). Since SFNO can be split across multiple GPUs during training, we can train large SFNOs and assess the effect of the SFNO size on the ensemble dispersion.
@@ 3 Ensemble Diagnostics
Ultimately, with SFNO-BVMC, we hope to analyze the statistics of LLHIs. This requires a calibrated ensemble with reliable probabilistic forecasts. SFNO-BVMC is a novel way to create ensemble forecasts from deterministic ML models. Therefore, in the following section, we present a diagnostics pipeline to evaluate the SFNO-BVMC ensemble and compare it to the IFS ensemble. We first evaluate SFNO-BVMC using diagnostics that evaluate overall performance. Next, we assess SFNO-BVMC’s control, perturbed, and ensemble mean spectra. Finally, we present diagnostics specifically focused on extreme weather forecasts. We open-source the code for these diagnostics (see Data Availability section), and we hope that it can be used to guide future ML model development. For a fair comparison for all diagnostics, we validate IFS against ECMWF’s operational analysis and SFNO-BVMC against ERA5. All diagnostics show SFNO-BVMC results with 58 members, and IFS ENS results with 50 members. SFNO-BVMC has 58 members: 29 checkpoints and 1 bred vector per checkpoint (added to and subtracted from the initial condition). Because of the use of 29 checkpoints and centered bred vector perturbations, SFNO-BVMC cannot be evaluated with an ensemble size smaller than fifty-eight members. However, while there are unbiased versions of the metrics, the difference in the metrics due to different ensemble size would be sufficiently small that the diagnostics still allow for fair evaluation.
@@ Acknowledgements.
The code, datasets, and models are all stored at https://doi.org/10.5061/dryad.2rbnzs80n. The code is integrated with Zenodo: https://doi.org/10.5281/zenodo.14710345. We include the code to train SFNO, conduct ensemble inference with bred vectors and multiple checkpoints, and scoring and analysis code. We also open-source the model weights of the trained SFNO. See the README of the former DOI for information on how to use the codebase and for the permissive license associated with the code and data.